# Viscous-inertial waves on the surface of the Sun: modeling, forward and inverse problems

### v2

@@ Abstract
This paper develops a mathematical framework for studying the newly discovered solar inertial oscillations, offering promising new avenues for exploring the Sun’s dynamics. Under the assumption of purely toroidal motions, the stream function of the flow satisfies a fourth-order scalar equation governing inertial waves on the rotating Sun. We prove well-posedness of wave solutions under explicit conditions on differential rotation. Moreover, we study the inverse problem of simultaneously reconstructing viscosity and differential rotation parameters from either complete or partial surface observations. To this end, we verify the tangential cone condition, ensuring the convergence of iterative regularization methods, furthermore the local unique identifiability of the unknown parameters. Numerical experiments employing the Nesterov-Landweber iteration confirm robustness of the reconstruction across different observation schemes and noise levels.
@@
inverse problems, differential equations on manifolds, forth order elliptic differential equations, helioseismology, inertial waves, iterative regularization, tangential cone condition
@@ Acknowledgement.
Support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1456/432680300 Mathematics of Experiment, Project C04 is gratefully acknowledged.
@@ 1 Introduction
Helioseismology is the study of the internal structure and dynamics of the Sun from observations of solar oscillations on the surface. Such inference is primarily done by analyzing the observed pressure wave (p-modes) and has led to many achievements including the determination of the Sun’s differential rotation [37]. However, the sensitivity of p-modes decreases with depth exacerbating the ill-posedness of the inversion. It is this ill-posedness that motivates the present article, and its accompanying analysis of regularized recovery of rotation and viscosity on the Sun.
The recent discovery of numerous inertial modes on the surface of the Sun [15] opens new possibilities for helioseismology. Inertial oscillations, whose frequencies are in the range of hundreds of nHz, are caused by the Coriolis force, exhibiting periodic behavior on the timescale of the solar rotation cycle. As a consequence, these modes require very long observation time to be identified. Indeed, their discovery relies on 10 years of data acquisition by the Helioseismic and Magnetic Imager on board the Solar Dynamics Observatory [34]. As many of these modes have maximum energy at the base of the convection zone [15], i.e., at a depth of about 0.71 solar radii, their study promises to give new insight into the dynamics of the deeper layers of the Sun. We will use these newly discovered waves as data of the inversion process.
Most of the observed modes are quasi-toroidal, i.e. almost divergence-free, with very small radial velocity component, giving rise to the use of a 2D model for inertial waves on a rotating sphere. In this framework, we can reduce the vectorial wave equation into a scalar equation employing a stream function, the divergence-free component of the Helmholtz decomposition of the flow. This idea was proposed in [40] and extended in [7] to take into account viscous stress, an important component in solar applications [12].
Inspired by these works, we will model the propagation of inertial-viscous waves on the surface of the rotating Sun by the stream function. The vectorial motion equation driven by Coriolis force and viscous stress will be turn into a fourth-order scalar equation in a low-dimensional setting. Moreover, we derive the separated equations for each longitudinal-wavenumber $m$ with realistic and physical meaningful $m$-dependent boundary conditions; this serves as the underlying model for the inverse parameter identification problem.
The problem of solar inversion using inertial waves is completely novel as these modes have been discovered only recently [15]. Here, we take a first step towards recovering the surface differential rotation and the viscosity from Dopplergrams. Doppler data are usually not used directly in helioseismology due to the high level of convective noise; indeed, solar excitations can be viewed as random turbulence, resulting in random oscillations. However, some inertial modes have been detected directly in the Doppler signal [14] motivating this study using direct measurements of inertial waves as input data.
Helioseismology predominantly solves linear inverse problem, under the assumption that perturbations are relatively small, as it is the case, for example, for meridional flows [13]. However, for parameters such as viscosity, where no suitable initial guess is available, this linearization assumption may no longer be valid. Here, following a related approach for pressure waves and correlation data in [27], we solve the nonlinear inversion for rotation and viscosity using inertial waves and iterative regularization reconstruction methods.
@@
This article provides a complete treatment – modeling, forward problem and inverse problem – for the newly observed inertial waves on the Sun:
After a rigorous derivation of the forward model already used in [12], we prove existence, uniqueness and stability of both, the full and separated the inertial wave equations. This well-posedness is derived under explicit conditions on the relative smallness of variations of the the latitudinal-dependent rotation $\Omega$ compared to the product of the frequency and the third power of the viscous parameter $\gamma$. In passing, these results also allow us to apply analytic Fredholm theory, providing mathematical foundations to the recent study of inertial modes. The results of our analysis, both on the conditions for uniqueness and on the isolated occurrence of inertial modes, agree with state-of-the-art interpretations in inertial mode helioseismology.
For the inverse problem of the simultaneous identification of viscosity and differential rotation, we lay out foundations of iterative regularization methods by characterizing adjoint operators on a continuous level. Crucially, we provide convergence guarantees for these methods – for full and, to a certain extent, also for restricted observations. We achieve this by establishing the important tangential cone condition via a lifted regularity strategy, accounting for realistic $L^{2}$-data. In addition, we prove local unique identifiability of $\Omega$ when $\gamma$ is known, and vice versa, from full measurement.
@@
Section 2 introduces the modeling framework for inertial oscillations and outlines different observation strategies for solar data. The well-posedness and regularity of the resulting wave solutions are derived in Section 3. Section 4 addresses the inverse problem for viscosity and differential rotation, including sensitivity analysis and derivation of associated adjoint operators. In Section 5, we establish convergence guarantees of iterative reconstruction schemes and local unique identifiability of the unknown parameters based on the tangential cone condition. Finally, Section 6 presents numerical reconstruction results under different scenarios, and Section 7 concludes our findings with a summary and an outlook on future research directions.
@@ 2.1 Reduced-order modeling of inertial waves
In this section, we derive a physical model for dynamics of waves in the presence of latitudinal differential rotation and eddy viscosity on the Sun in a rotating frame.
@@
We begin with the equation of motion (momentum equation), in which the dynamics of a moving particle of density $\rho$ in an incompressible fluid is subject to linear viscous stress $\mathbf{\bm{\tau}}$ with viscosity $\gamma$, acoustic pressure $p$, gravity $\mathbf{g}$, and external source $\mathbf{f}$:
@@
Here, $\textbf{{v}}=\textbf{{v}}(\textbf{{r}},t)\in\mathbb{R}^{3}$ is the vector velocity of a particle at position $\textbf{{r}}\in\mathbb{R}^{3}$ and time $t\in(0,T)$, while $\nabla_{\!\textbf{{v}}}:=\sum_{i=1}^{3}v_{i}\frac{\partial}{\partial x_{i}}$ applies componentwise. We work in the frame rotating at the angular frequency $\Omega_{\mathrm{{ref}}}$ around a fixed axis $\mathbf{\hat{a}}$ and set $\mathbf{\Omega}_{\mathrm{ref}}=\Omega_{\mathrm{{ref}}}\,\mathbf{\hat{a}}$. Equation (1a) corresponds to the Navier-Stokes equation in a rotating frame. The fictitious force $(2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{v}})$ is the inertial Coriolis force resulting from the transformation between the fixed and the rotating frame. The other additional fictitious forces, Euler force $\rho\left(\frac{d\mathbf{\Omega}_{\mathrm{ref}}}{dt}\times\textbf{{r}}\right)$ and centrifugal force $\rho\mathbf{\Omega}_{\mathrm{ref}}\times(\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{r}})$, are ignored as $\Omega_{\mathrm{{ref}}}$ is uniform in time and assumed to be small. For solar applications, one usually chooses the Carrington frame as reference frame with the angular velocity $\Omega_{\mathrm{{ref}}}=14.7^{0}/$day and the rotation axis $\mathbf{e}_{z}=[0,0,1]^{\top}$. Equation (1a) is complemented by the continuity equation (conservation of mass):
@@
$\textbf{{v}}(\textbf{{r}},t)={\textbf{{u}}_{0}}(\textbf{{r}})+\textbf{{u}}(\textbf{{r}},t),\quad\rho(\textbf{{r}},t)=\rho_{0}(\textbf{{r}})+\rho^{\prime}(\textbf{{r}},t),$
$p(\textbf{{r}},t)=p_{0}(\textbf{{r}})+p^{\prime}(\textbf{{r}},t),\quad\mathbf{g}(\textbf{{r}},t)=\mathbf{g}_{0}(\textbf{{r}})+\mathbf{g}^{\prime}(\textbf{{r}},t),$
where u is the (perturbed) wave velocity, and ${\textbf{{u}}_{0}}(\textbf{{r}}):=(\mathbf{\Omega}(\textbf{{r}})-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{r}}$ is assumed to be a differential rotation field with angular velocity $\mathbf{\Omega}(\textbf{{r}})=\Omega(r,\theta)\mathbf{e}_{z}$ depending only on radius $r$ and colatitude $\theta$ in spherical coordinates $r;\theta;\phi]\in[0,\infty)\times[0,\pi)\times[0,2\pi)$. We assume that the background medium solves the second-order quasilinear elliptic boundary value problem
@@
The background gravity force derives from a potential $\mathbf{g}_{0}=-\nabla\Phi_{0}$ that satisfies the Poisson equation $\Delta\Phi_{0}=4\pi G\rho_{0}$, where $G$ is the gravitational constant. To construct such a background model for the Sun, it is usually assumed that the background flows are weak and the background medium is built from the hydrostatic equilibrium, that is $\operatorname{\mathbf{grad}}p_{0}=\rho_{0}\mathbf{g}_{0}$. This assumption leads to background coefficients that depend only on depth and are given by standard solar models such as model S [8]. Starting from a radial model for density, it is possible to solve Poisson equation to obtain the gravitational potential and deduce the background pressure.
@@
Waves have small amplitude in comparison to the mean flow, i.e. $|\textbf{{u}}|\ll|{\textbf{{u}}_{0}}|$. Moreover, the anelastic approximation $\rho^{\prime}\approx 0$ and the Cowling approximation $\mathbf{g}^{\prime}\approx 0$ (see [9]) hold true.
The anelastic approximation is justified if the flow speed is significantly smaller than the sound speed, which is adequate for inertial modes – a few meters per second versus kilometers per second for the surface sound speed. The system (1) thus becomes
$\rho_{0}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)=-\operatorname{\mathbf{grad}}p^{\prime}+\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\rho_{0}\mathbf{f},$ (3a)
@@
where $\bm{\tau}^{\prime}:=\operatorname{\mathbf{grad}}\textbf{{u}}+(\operatorname{\mathbf{grad}}\textbf{{u}})^{\top}$.
The flow field $\textbf{{u}}_{0}$}, equivalently $\Omega$, and the viscosity $\gamma$ will be the quantities of interest in the inverse problem.
Stratified equations.
We first rewrite the left-hand side of (3a) in spherical coordinates, i.e. in terms of the components $u_{r},u_{\theta},u_{\phi}$ of the expansion $\textbf{{u}}=u_{r}\mathbf{e}_{r}+u_{\theta}\mathbf{e}_{\theta}´+u_{\phi}\mathbf{e}_{\phi}$ into the local orthonormal basis vectors $\mathbf{e}_{r},\mathbf{e}_{\theta}$, and $\mathbf{e}_{\phi}$ as defined in the appendix A. Noting that ${\textbf{{u}}_{0}}=(\Omega(r,\theta)-\Omega_{\mathrm{{ref}}})\sin\theta\mathbf{e}_{\phi}$ and using the identities
@@
where $\mathbf{D}_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ is the material derivative.
We now introduce the $3\times 2$-matrix $\mathbf{R}(\textbf{{r}}):=[\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})]$. Then multiplication by the $3\times 3$ matrix $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ is the projection onto the tangential subspace and equivalent to the application of $\textbf{{r}}\times\textbf{{r}}\times$. By applying $\textbf{{r}}\times\textbf{{r}}\times$ to (3a), we wish to obtain an approximate equation for the horizontal velocity components $\textbf{{u}}_{\mathrm{h}}(\textbf{{r}}):=-\textbf{{r}}\times\textbf{{r}}\times\textbf{{u}}(\textbf{{r}})$ given by $\textbf{{u}}_{\mathrm{h}}=\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}\textbf{{u}}=u_{\theta}\mathbf{e}_{\theta}+u_{\phi}\mathbf{e}_{\phi}$. To this end, we need the following assumption:
@@
Due to this assumption, we can neglect the first column of the matrix in Eq. (4), and the application of $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ “kills” the first row of this matrix (the first element of the vector in (4) may be balanced by $\frac{\partial p^{\prime}}{\partial r}$). Therefore, we are left with the matrix
@@
and obtain the following system of equations for the tangential vector field $\textbf{{u}}_{\mathrm{h}}$ on $r{\mathbb{S}^{2}}$} for a fixed $r>0$ with right hand side $\textbf{{f}}_{\!\mathrm{h}}:=\mathbf{R}\mathbf{R}^{\top}\textbf{{f}}$:
@@
employing the fact $\operatorname{\mathbf{grad}}(\rho_{0}(\operatorname{\mathbf{grad}}\textbf{{u}}_{\mathrm{h}})^{\top})=\operatorname{\mathbf{grad}_{\mathrm{h}}}(\rho_{0}\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}})=0$ due to (5b).
Physically, the assumption of small radial motions is motivated by an identification of the observed inertial modes from [15] as the eigenvalues of the 3D linearized equations of momentum, mass and energy [5]. In this setup, the radial velocities of the eigenfunctions are small compared to the horizontal ones for most of the observed modes. This is also the case in nonlinear simulations of rotating convection in a spherical shell [6]. This assumption could possibly be justify mathematically within an asymptotic analysis as $r\to\infty$ or $\Omega\to 0$. This, however, would require to amend the system (3) by an atmospheric model and assumptions on the coefficient functions and is beyond the scope of this paper.
@@
$\rho_{0}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi.$
In order to obtain an equation for $\Psi$, we apply $\operatorname{curl_{\mathrm{h}}}$ to (5a). Noting that $\operatorname{curl_{\mathrm{h}}}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi=-\Delta_{\mathrm{h}}\Psi$ and using the identities
@@
$\operatorname{curl_{\mathrm{h}}}\mathbf{M}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}+\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},\quad\alpha_{\Omega}(\theta):=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\right)$

### v3

# Linear toroidal-inertial waves on a differentially rotating sphere with application to helioseismology: Modeling, forward and inverse problems

@@ Abstract
This paper develops a mathematical framework for interpreting observations of solar inertial waves in an idealized setting. Under the assumption of purely toroidal linear waves on the sphere, the stream function of the flow satisfies a fourth-order scalar equation. We prove well-posedness of wave solutions under explicit conditions on differential rotation. Moreover, we study the inverse problem of simultaneously reconstructing viscosity and differential rotation parameters from either complete or partial surface data. We establish convergence guarantee of iterative regularization methods by verifying the tangential cone condition, and prove local unique identifiability of the unknown parameters. Numerical experiments with Nesterov-Landweber iteration confirm reconstruction robustness across different observation strategies and noise levels.
@@
Inverse problems, differential equations on manifolds, fourth-order elliptic differential equations, helioseismology, inertial waves, iterative regularization, tangential cone condition.
@@ Acknowledgement.
Support from Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through SFB 1456/432680300 Mathematics of Experiment, Project C04 is gratefully acknowledged. LG and DF acknowledge funding from the ERC Synergy Grant WHOLESUN 810218.
@@ 1 Introduction
Helioseismology is the study of the internal structure and dynamics of the Sun from observations of solar oscillations on the surface. Such inference is primarily done by analyzing the observed pressure waves (p-modes) with periods of order five minutes. Acoustic-mode helioseismology based on the interpretation of the mode frequencies has led to many achievements, including the determination of the Sun’s differential rotation as a function of radius and unsigned latitude [Thompson03]. Inversions of the two-point correlations of the wave field at the surface can also be used to infer 3D perturbations in sound speed and flows in the solar interior [GizonBirchReview, Bjoern23].
The recent discovery of global Rossby modes on the Sun [Loptien2018] has opened a new research topic: inertial-mode helioseismology. Solar inertial oscillations are modes restored by the Coriolis force, with periods spanning several weeks, comparable to solar rotation period of $\sim 27$ days at the equator. These modes require very long observations from space-based and ground-based observatories to be studied with sufficient frequency resolution; 50 years of data are now available [Gizon2024, Liang2025]. Since many of these modes have maximum kinetic energy deep in the convective envelope of the Sun [GizonInertial21], their study promises to reveal new insights into the physics and dynamics of the solar interior. Current efforts focus on the development of simplified physical models of these inertial modes to investigate their unique sensitivity to internal properties of the Sun, such as differential rotation at very high latitudes and the internal turbulent viscosity. Several highly idealized 2D models focusing on purely toroidal modes have been developed for both $\beta$-plane [Gizon2020] and spherical geometries [Damien22]. The dynamics of these retrograde-propagating modes is governed by the functional form of the background latitudinal differential rotation, with slower rotation at high latitudes than at the equator. Particularly, 2D models have demonstrated the existence of viscous critical latitudes for most inertial modes.
In this article, by employing stream function formulation in a 2D spherical framework, we reduce the vectorial viscous-inertial wave equation to a fourth-order scalar equation of Orr-Sommerfeld type [Gizon2020, Damien22]. Additionally, we derive the separated equations for each longitudinal-wavenumber $m$ with realistic and physically meaningful $m$-dependent boundary conditions at the poles. The separated equations serve as the underlying models for our inverse problem: the retrieval of differential rotation and turbulent viscosity. Since solar inertial wave inversion is completely novel in helioseismology, this work establishes the first step towards extracting these quantities from observed surface horizontal velocities.
@@
This article provides a comprehensive treatment – modeling, forward problem and inverse problem – of purely toroidal inertial modes governed by a time harmonic fourth-order equation on a differentially rotating sphere.
After a derivation of the forward model used in [Damien22], we prove existence, uniqueness and stability of full and separated inertial wave equations. Well-posedness is guaranteed under explicit conditions that the latitudinal-dependent rotation $\Omega$ is small relative to the product of the frequency and the third power of the viscousity $\gamma$. These results enable the application of analytic Fredholm theory, providing a mathematical foundation for the model [Damien22], and rigorously validating structure of discrete isolated inertial modes.
We develop a regularization framework for the inverse problem of simultaneously reconstructing viscosity and differential rotation. The framework is built on characterization of continuous adjoint operators and convergence guarantees for iterative regularization methods – a result we obtain for both full and, under certain condition, for restricted observations. Central to our convergence analysis is the establishment of the tangential cone condition through a lifted regularity strategy, accommodating realistic $L^{2}$-data. Furthermore, we prove local unique identifiability: $\Omega$ is uniquely identifiable when $\gamma$ is known, and vice versa under full measurement.
@@
The article is organized as follows. Section 2 introduces the modeling framework for inertial oscillations and outlines different observation strategies for solar data. Section 3 establish well-posedness and regularity of the resulting wave solutions. Section 4 addresses the inverse problem for viscosity and differential rotation, including sensitivity analysis and adjoint operator derivation. In Section 5, we develop convergence guarantees for iterative reconstruction methods and prove local unique identifiability via the tangential cone condition. Section 6 demonstrate numerical reconstruction performance across different measurement scenarios. Finally, Section 7 concludes our findings and future directions.
@@ 2 Model and observation of inertial waves
We begin with deriving a model for dynamics of linear inertial waves on the Sun in a rotating frame, incorporating latitudinal differential rotation and eddy viscosity. The theoretical framework is developed in Sections 2.12.2, while observations of solar data are discussed in Section 2.3.
@@ 2.1 Reduced-order modeling of inertial waves
We begin with the equation of motion (momentum equation), in which the dynamics of a moving particle of density $\rho$ is subject to linear viscous stress $\mathbf{\bm{\tau}}$ with viscosity $\gamma$, acoustic pressure $p$, gravity $\mathbf{g}$, and external source $\mathbf{f}$:
@@
Here, $\textbf{{v}}=\textbf{{v}}(\textbf{{r}},t)\in\mathbb{R}^{3}$ is the vector velocity of a particle at position $\textbf{{r}}\in\mathbb{R}^{3}$ and time $t\in(0,T)$, while $\nabla_{\!\textbf{{v}}}:=\sum_{i=1}^{3}v_{i}\frac{\partial}{\partial x_{i}}$ applies componentwise. We work in the frame rotating at the constant angular frequency $\Omega_{\mathrm{{ref}}}$ around a fixed axis $\mathbf{\hat{a}}$ and set $\mathbf{\Omega}_{\mathrm{ref}}=\Omega_{\mathrm{{ref}}}\,\mathbf{\hat{a}}$. Equation (1a) corresponds to the Navier-Stokes equation in a rotating frame. The fictitious force $(2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{v}})$ is the inertial Coriolis force resulting from the transformation between the fixed and the rotating frame. The other additional fictitious forces, Euler force $\rho\left(\frac{d\mathbf{\Omega}_{\mathrm{ref}}}{dt}\times\textbf{{r}}\right)$ and centrifugal force $\rho\mathbf{\Omega}_{\mathrm{ref}}\times(\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{r}})$, are ignored as $\Omega_{\mathrm{{ref}}}$ is uniform in time and assumed to be small. For solar applications, one usually chooses the Carrington frame as reference frame with the angular velocity $\Omega_{\mathrm{{ref}}}=14.7$ deg/day and the rotation axis $\mathbf{e}_{z}=[0,0,1]^{\top}$. Equation (1a) is complemented by the continuity equation (conservation of mass):
@@
$\textbf{{v}}(\textbf{{r}},t)={\textbf{{u}}_{0}}(\textbf{{r}})+\textbf{{u}}(\textbf{{r}},t),\qquad\rho(\textbf{{r}},t)=\rho_{0}(\textbf{{r}})+\rho^{\prime}(\textbf{{r}},t),$
$p(\textbf{{r}},t)=p_{0}(\textbf{{r}})+p^{\prime}(\textbf{{r}},t),\qquad\mathbf{g}(\textbf{{r}},t)=\mathbf{g}_{0}(\textbf{{r}})+\mathbf{g}^{\prime}(\textbf{{r}},t),$
where u is the (perturbed) wave velocity, and ${\textbf{{u}}_{0}}(\textbf{{r}}):=(\mathbf{\Omega}(\textbf{{r}})-\mathbf{\Omega}_{\mathrm{ref}})\times\textbf{{r}}$ is assumed to be a differential rotation field with angular velocity $\mathbf{\Omega}(\textbf{{r}})=\Omega(r,\theta)\mathbf{e}_{z}$ depending only on radius $r$ and colatitude $\theta$ in spherical coordinates $r;\theta;\phi]\in[0,\infty)\times[0,\pi)\times[0,2\pi)$. This linearization is possible when no linearly unstable mode exists. In this regime, there is no growing (in time) solutions to Eq. (1). Note that such solutions exists in nonlinear simulations with solar-like differential rotation [Bekki2024], however this regime is not included in this study.
Background medium.
We assume that the background medium solves the second-order quasilinear elliptic boundary value problem
@@
The background gravity force derives from a potential $\mathbf{g}_{0}=-\nabla\Phi_{0}$ that satisfies the Poisson equation $\Delta\Phi_{0}=4\pi G\rho_{0}$, where $G$ is the gravitational constant. To construct such a background model for the Sun, it is usually assumed that the background flows are weak and the background medium is built from the hydrostatic equilibrium, that is $\operatorname{\mathbf{grad}}p_{0}=\rho_{0}\mathbf{g}_{0}$. This assumption leads to background coefficients that depend only on depth and are given by standard solar models such as Model S [modelS]. Starting from a radial model for density, it is possible to solve Poisson equation to obtain the gravitational potential and deduce the background pressure.
First-order equations.
The first-order the system (1) then becomes
$\rho_{0}\left(\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}\right)=-\operatorname{\mathbf{grad}}p^{\prime}+\rho^{\prime}\mathbf{g}_{0}+\rho_{0}\mathbf{g}^{\prime}+\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\rho_{0}\mathbf{f},$ (3a)
$\frac{\partial\rho^{\prime}}{\partial t}+\mathrm{div}(\rho^{\prime}\mathbf{u}_{0})+\mathrm{div}(\rho_{0}\mathbf{u})=0,$ (3b)
where $\bm{\tau}^{\prime}:=\operatorname{\mathbf{grad}}\textbf{{u}}+(\operatorname{\mathbf{grad}}\textbf{{u}})^{\top}$. It is complemented by the linearized equation of state
$\frac{p^{\prime}}{p_{0}}=\gamma\frac{\rho^{\prime}}{\rho_{0}}+\frac{s^{\prime}}{c_{v}},$
where $\gamma=5/3$ is the ratio of specific heats, the quantity $c_{v}$ is the specific heat at constant volume and $s^{\prime}$ denotes the perturbed entropy. The first two terms of the right hand side of Eq. 3a can be rewritten as
$\operatorname{\mathbf{grad}}p^{\prime}+\rho^{\prime}\mathbf{g}_{0}=-\rho_{0}\operatorname{\mathbf{grad}}\left(\frac{p^{\prime}}{\rho_{0}}\right)+p^{\prime}\frac{N^{2}}{\mathbf{g}},$ (4)
where we used hydrostatic equilibrium and defined the buoyancy frequency $N$ such that
$N^{2}=\mathbf{g}\cdot\left(-\frac{\nabla\rho_{0}}{\rho_{0}}+\frac{\nabla p_{0}}{\gamma p_{0}}\right).$ (5)
@@
We employ the Cowling approximation with $\mathbf{g}^{\prime}=0$ (see [Cowling1941]), which is classic in helioseismology. We further assume the anelastic approximation which is justified if the flow speed is significantly smaller than the sound speed. This is adequate for inertial modes – flow speeds of a few meters per second compared to kilometers per second for the surface sound speed – and has been numerically verified [Sup2025]. Two additional assumptions enable reduction to a 2D problem on the sphere: adiabaticity ($s^{\prime}=0$), and strongly stratified medium (see also [Watson81]) so that the term in $N^{2}$ in Eq. 4 can be neglected.
Under these assumptions, the system (3) becomes
$\frac{\partial\textbf{{u}}}{\partial t}+\nabla_{\!{\textbf{{u}}_{0}}}\textbf{{u}}+\nabla_{\!\textbf{{u}}}{\textbf{{u}}_{0}}+2\mathbf{\Omega}_{\mathrm{ref}}\times\textbf{{u}}=-\operatorname{\mathbf{grad}}\left(\frac{p^{\prime}}{\rho_{0}}\right)+\frac{1}{\rho_{0}}\operatorname{\mathbf{div}}\left[\rho_{0}\gamma\bm{\tau}^{\prime}\right]+\mathbf{f},$ (6a)
@@
where the flow field $\textbf{{u}}_{0}$}, equivalently $\Omega$, and the viscosity $\gamma$ will be the quantities of interest in the inverse problem.
First-order equations on the sphere.
We now rewrite the left hand side of (6a) in spherical coordinates using the components $u_{r},u_{\theta},u_{\phi}$ in the expansion $\textbf{{u}}=u_{r}\mathbf{e}_{r}+u_{\theta}\mathbf{e}_{\theta}´+u_{\phi}\mathbf{e}_{\phi}$, where $\mathbf{e}_{r}$, $\mathbf{e}_{\theta}$, $\mathbf{e}_{\phi}$ are the local orthonormal basis vectors defined in the appendix A. Noting that ${\textbf{{u}}_{0}}=(\Omega(r,\theta)-\Omega_{\mathrm{{ref}}})\sin\theta\mathbf{e}_{\phi}$ and using the identities
@@
where $\mathbf{D}_{t}:=\frac{\partial}{\partial t}+(\Omega-\Omega_{\mathrm{{ref}}})\frac{\partial}{\partial\phi}$ is the material derivative acting on vector fields.
Define the $3\times 2$-matrix $\mathbf{R}(\textbf{{r}}):=[\mathbf{e}_{\theta}(\textbf{{r}}),\mathbf{e}_{\phi}(\textbf{{r}})]$. Then multiplication by the $3\times 3$ matrix $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ is the projection onto the tangential subspace and equivalent to the application of $\textbf{{r}}\times\textbf{{r}}\times$. By applying $\textbf{{r}}\times\textbf{{r}}\times$ to (6a), we wish to obtain an approximate equation for the horizontal velocity components $\textbf{{u}}_{\mathrm{h}}(\textbf{{r}}):=-\textbf{{r}}\times\textbf{{r}}\times\textbf{{u}}(\textbf{{r}})$ given by $\textbf{{u}}_{\mathrm{h}}=\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}\textbf{{u}}=u_{\theta}\mathbf{e}_{\theta}+u_{\phi}\mathbf{e}_{\phi}$. To this end, we require the following assumption:
@@
Under this assumption, one can neglect the first column of the matrix in Eq. (7), and the application of $\mathbf{R}(\textbf{{r}})\mathbf{R}(\textbf{{r}})^{\top}$ “kills” the first row of this matrix (the first element of the vector in (7) may be balanced by $\frac{\partial p^{\prime}}{\partial r}$). Therefore, we are left with the matrix
@@
and obtain the following system of equations for the tangential vector field $\textbf{{u}}_{\mathrm{h}}$ on the sphere $r{\mathbb{S}^{2}}$} for a fixed radius $r>0$ as:
@@
with the right hand side $\textbf{{f}}_{\!\mathrm{h}}:=\mathbf{R}\mathbf{R}^{\top}\textbf{{f}}$. Here, we employ the fact $\operatorname{\mathbf{grad}}(\rho_{0}(\operatorname{\mathbf{grad}}\textbf{{u}}_{\mathrm{h}})^{\top})=\operatorname{\mathbf{grad}_{\mathrm{h}}}(\rho_{0}\operatorname{div_{\mathrm{h}}}\textbf{{u}}_{\mathrm{h}})=0$ due to (8b), with operators $\bm{\Delta}_{\mathrm{h}}$, $\operatorname{\mathbf{grad}_{\mathrm{h}}}$, $\operatorname{div_{\mathrm{h}}}$ defined in (51)-(52) in Appendix A. We note that all differential operators with subscript $\rm h$} denote operators on spheres and are defined in Appendix A.
Physically, the assumption of small radial motions is motivated by an identification made in [GizonInertial21], namely, that the observed inertial modes can be viewed as eigenvalues of the 3D linearized equations of momentum, mass and energy. In this setting, the radial velocities of the eigenfunctions are small relative to the horizontal components for most observed modes. This observation is also supported by nonlinear simulations of rotating convection in a spherical shell [Bekki2022]. While this assumption could potentially be justified through asymptotic analysis ($r\to\infty$ or $\Omega\to 0$), such a rigorous treatment would require amending the system (6) with an atmospheric model and explicit assumptions on the coefficient functions – a task beyond this paper’s scope.
@@
$\rho_{0}\textbf{{u}}_{\mathrm{h}}=\operatorname{\mathbf{curl}_{\mathrm{h}}}(\Psi)$
with $\operatorname{\mathbf{curl}_{\mathrm{h}}}$ as in (51)-(52). In order to obtain an equation for $\Psi$, we apply $\operatorname{curl_{\mathrm{h}}}$ to (8a). Noting that $\operatorname{curl_{\mathrm{h}}}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\operatorname{curl_{\mathrm{h}}}\operatorname{\mathbf{curl}_{\mathrm{h}}}\Psi=-\Delta_{\mathrm{h}}\Psi$ and using the identities
@@
$\operatorname{curl_{\mathrm{h}}}\mathbf{M}(\rho_{0}\textbf{{u}}_{\mathrm{h}})=\alpha_{\Omega}\frac{\partial\Psi}{\partial\phi}+\frac{1}{r^{2}}\frac{\partial\Omega}{\partial\theta}\frac{\partial^{2}\Psi}{\partial\theta\,\partial\phi},\qquad\alpha_{\Omega}(\theta):=\frac{1}{r^{2}\sin\theta}\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d}{d\theta}(\Omega(\theta)\sin^{2}\theta)\right)$